Pina Colada Corp.’s balance sheet at December 31, 2018, is presented below.

Pina Colada Corp.
Balance Sheet
December 31, 2018

Cash $14,300 Accounts payable $8,900
Accounts receivable 20,600 Common stock 19,500
Allowance for doubtful accounts (700 ) Retained earnings 14,900
Inventory 9,100
$43,300 $43,300


During January 2019, the following transactions occurred. Pina Colada uses the perpetual inventory method.

Jan. 1 Pina Colada accepted a 4-month, 8% note from Merando Company in payment of Merando’s $5,100 account.
3 Pina Colada wrote off as uncollectible the accounts of Inwood Corporation ($500) and Goza Company ($100).
8 Pina Colada purchased $16,700 of inventory on account.
11 Pina Colada sold for $26,600 on account inventory that cost $17,600.
15 Pina Colada sold inventory that cost $800 to Mark Lauber for $1,100. Lauber charged this amount on his Visa First Bank card. The service fee charged Pina Colada by First Bank is 3%.
17 Pina Colada collected $20,200 from customers on account.
21 Pina Colada paid $16,400 on accounts payable.
24 Pina Colada received payment in full ($100) from Goza Company on the account written off on January 3.
27 Pina Colada purchased supplies for $1,400 cash.
31 Pina Colada paid other operating expenses, $3,500.


Adjustment data:

1. Interest is recorded for the month on the note from January 1.
2. Bad debts are expected to be 6% of the January 31, 2019, accounts receivable.
3. A count of supplies on January 31, 2019, reveals that $500 remains unused.

Prepare an adjusted trial balance at January 31, 2019. (Round answers to 0 decimal places, e.g. 5,275.)

Pina Colada Corp.
Adjusted Trial Balance

For the Quarter Ended January 31, 2019January 31, 2019For the Year Ended January 31, 2019

Debit

Credit

$

$

Total $

$
Answer #1
Trial balance
Account Titles Debit Credit
Cash $          13,267 =14300-1100*3%+20200-16400+100-1400-3500
Accounts Receivable $          22,400 =20600-5100-600+1100+26600-20200+100-100
Allowance for doubtful accounts $             1,344 =700-600+100+1144
Note Receivable $            5,100 =5100
Interest Receivable $                  34 =5100*8%*1/12
Inventory $            7,400 =9100+16700-800-17600
Supplies $                500 =1400-900
Accounts Payable $             9,200 =8900+16700-16400
Common Stock $           19,500 =19500
Retained Earnings $           14,900 =14900
Sales Revenue $           27,700 =1100+26600
Interest Revenue $                   34 =5100*8%*1/12
Cost of Goods Sold $          18,400 =800+17600
Service Fee $                  33 =1100*3%
Operating Expenses $            3,500 3500
Bad Debt expense $            1,144 =22400*6%-200
Supplies Expense $                900 =1400-500
Total $          72,678 $           72,678